OpenStudy (anonymous):

Is (2, 3) a solution to this system of equations? 8x + 7y = 37   2y = 4x - 2

OpenStudy (anonymous):

Im not entirely sure but I believe you plug 2 in for x and 3 in for y within the original equation..If you solve and the statement is correct then it is a solution

OpenStudy (anonymous):

Well, there's three ways we could solve this. We could solve by elimination, substitution, or by graphing. Here's a graph: is the given point a solution?

OpenStudy (anonymous):

I would use the elimination process, which is: 8x+7y=37 +-4x+2y=-2 The easiest thing to do is eliminate the x first , so you need to multiply the secon equation by 2 .. so the new system of equations is: 8x+7y=37 + -8x+4y=-4 The x cancels out and you're left with : 11x=33 ... Divide 11 on both sides and you get x=3 , now plug 3 into one of the equations for x to find y.